OpenStudy (anonymous):

Identify which property of equality is used to transform Equation 1 to Equation 2. Equation 1: x – 6 = 14 Equation 2: x = 20 A. addition property of equality B. subtraction property of equality C. multiplication property of equality D. division property of equality Plz Explain

OpenStudy (knov):

x=14+6

OpenStudy (anonymous):

I think it is A

OpenStudy (knov):

to find out x you need to bring the all values behind the = which means in this equation x=14+6. Maybe you wonder why there is + instead of -. bcuz when symbols taken behind the = they go inverse. for exampe x-2=5 is x=5+2, x+8=9 is x=9-8

OpenStudy (knov):

yup it's a addition of course.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (knov):

no problem